# Matrix systems, algebras, and open maps

### v1

@@ Abstract.
open map, numerical range, matrix system, quantum state
@@ 1. Introduction
So $\mathcal{O}|_{\mathcal{R}}:=\{\ell|_{\mathcal{R}}:\ell\in\mathcal{O}\}$ contains $\omega_{\lambda}|_{\mathcal{R}}$ but none of the restriction $\ell_{\theta}|_{\mathcal{R}}$ of
@@
Corey et al. [18] and Leake et al. [29, 30] first studied a problem of numerical ranges closely related to (1.1). The problem (1.1) was studied by Weis [52, 53] and Rodman et al. [46] when $\mathcal{D}(\mathcal{R})$ is replaced with the affinely isomorphic joint numerical range. Numerical ranges are the topic of Sec. 4.
@@
Thm. 1.2 can simplify the problem (1.1), and related continuity problems, if $\mathcal{R}$ is included in a *-subalgebra of $\mathrm{M}_{n}$ smaller than $\mathrm{M}_{n}$, as we show in Sec. 8. Examples from quantum information theory are presented in Sec. 9.
@@ 2. States and density matrices
The aim of this section is to translate openness questions from states to density matrices. In the sequel, we refer to matrix systems and *-subalgebras synonymously as complex matrix systems and complex *-subalgebras, respectively. We introduce their real counterparts because real *-subalgebras have a greater variety of state spaces than the complex ones.
@@
Real and complex matrix systems have the same families of state spaces. If $\mathcal{R}$ is a real or complex matrix system on $\mathbb{C}^{n}$, then the real matrix system of hermitian matrices $\mathcal{H}(\mathcal{R})$ has the same set of hermitian matrices and hence the same set of density matrices as $\mathcal{R}$. Conversely, any real matrix system $\mathcal{R}$ included in $\mathcal{H}(\mathrm{M}_{n})$ has the same same set of hermitian matrices and the same set of density matrices as the complex matrix system $\mathcal{R}\oplus\mathrm{i}\mathcal{R}$. The set of density matrices of a real or complex matrix system $\mathcal{R}$ is affinely isomorphic the state space of $\mathcal{R}$ by Lemma 2.1 or 2.2, respectively.
@@
Indeed, $\mathcal{C}_{2}^{\vee}=\mathcal{C}_{2}$ holds [28, Thm. III.2.1] as the space of hermitian matrices $\mathcal{H}_{2}$ is a Euclidean Jordan algebra with Jordan product $A\circ B=\frac{1}{2}(AB+BA)$ and inner product $(A,B)\mapsto\mathrm{Re}\langle A,B\rangle$, $A,B\in\mathcal{H}_{2}$. So, $\mathcal{C}_{2}^{\vee}=\mathcal{C}_{2}\subset\mathcal{C}_{1}\subset\mathcal{C}_{1}^{\vee}$ proves $\mathcal{C}_{2}^{\vee}=\mathcal{C}_{2}$, which implies (2.4). See also the Notes to Chapter 6 in [2].
@@ 3. Direct convex sums
As $K:=K_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}K_{m}$ is a metric space, it suffices to show that there are arbitrary small family members and $O:=O_{I}(\epsilon,(A_{i})_{i\in I})$ is open.
@@
    Thereby, the equivalence of states and density matrices is described in Coro. 2.7. It is important to observe that $\frac{1}{2}(\mathds{1}_{2}+\cos(\theta)X+\sin(\theta)Z)\oplus 0$ and $\cos(\theta)(X\oplus 1)+\sin(\theta)Z\oplus 0$ are matrices in $\mathrm{M}_{2}\oplus\mathrm{M}_{1}$ for all $\theta\in\mathbb{R}$. The map $\pi$ is open at every point in the complement of $\mathcal{G}_{0}$ because $\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})$ is a cone over a ball, see [52, Lemma 4.17] for a detailed proof.
@@
over $M$. The map $\pi$ is not open at any point in the half-open segment $\mathcal{G}_{0}$ and open at every point in the complement. The lack of openness can be visualized graphically in three-space by observing that $\pi$ projects the cone $\mathcal{D}(\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R}))$ along its generatrix $\mathcal{G}$ to the ellipse $\mathcal{D}(\mathcal{R})$.
@@ 4. Numerical ranges
$f_{A}:\mathbb{C}S^{n}\to W(A)\,,\quad\ket{\varphi}\mapsto\braket{\varphi}{A\varphi}\,.$ (4.2)
@@
If $z\in W(A)$ is not an extreme point of $W(A)$, then $f_{A}^{-1}$ is strongly continuous at $z$ [18, Thm. 4]. Let $z$ be an extreme point of $W(A)$. Then there are $\theta_{0}\in\mathbb{R}$ and $i_{0}\in\{1,\ldots,n\}$ such that $z=z_{i_{0}}(\theta_{0})$. Now, $f_{A}^{-1}$ is strongly continuous at $z$ if and only if $z_{i}(\theta_{0})=z$ implies $z_{i}=z_{i_{0}}$ for all $i=1,\ldots,n$, see [30, Thm. 2.1.1]. Leake at al. [30] state the latter condition by saying that the eigenfunctions corresponding to $z$ at $\theta_{0}$ do not split.
@@
Strong continuity is connected to the openness of a linear map. For all $z\in W(A)$, the multi-valued map $f_{A}^{-1}$ is strongly continuous at $z$ if and only if the restricted linear map $v:\mathcal{D}(\mathrm{M}_{n})\to V(A_{1},A_{2})$, introduced in (4.1), is open at every point in the fiber $v^{-1}[(\mathrm{Re}(z),\mathrm{Im}(z)){}^{\mathrm{T}}]$, see [53, Coro. 5.2]. Moreover, for every $x\in V(A_{1},A_{2})$, the map $v$ is open at some point in the relative interior [^2] The relative interior [45] of a convex set $K$, denoted by $\operatorname{ri}(K)$, is the interior of $K$ in the relative topology of the affine hull of $K$. of $v^{-1}(x)$ if and only if $v$ is open at every point in $v^{-1}(x)$.
@@ 5. Retractions
    The stability problem of a finite-dimensional compact convex set $K$ is completely solved [40]. The $d$-skeleton of $K$ is the union of all faces [^3] A face [45] of a convex set $K$ is a convex subset $F\subset K$ such that $x,y\in F$ is implied by $(1-s)x+sy\in F$ for all $x,y\in K$ and $s\in(0,1)$. of $K$ of dimension at most $d$. The convex set $K$ is stable if and only if for every nonnegative integer $d$, the $d$-skeleton of $K$ is closed.
@@
    The set of density matrices $\mathcal{D}(\mathrm{M}_{n})$ is stable [20] because all its $d$-skeletons are closed. The closedness follows from three arguments: First, the set $\mathcal{D}(\mathrm{M}_{n})$ is a compact convex set of dimension $n^{2}-1$. Second, every nonempty face of $\mathrm{M}_{n}$ is unitarily similar to $\mathcal{D}(\mathrm{M}_{l})\oplus\{0\}$ for some positive integer $l$, and third, and the unitary group $U(n)$ is compact.
@@
$K^{\times k}\to K\,,\quad(x_{1},\dots,x_{k})\mapsto\frac{1}{k}\sum_{i=1}^{k}x_{i}$
the $k$-fold cartesian product of $K$ is open for all positive integers $k$. More generally, for every tuple $(s_{1},\dots,s_{k})$ of nonnegative real numbers adding up to one, the map
@@ 8. Topology simplified by algebra
As a word of explication, the map $\pi$ factors by Lemma 2.5. The claim follows from Lemma 8.1 a), by letting $\mathcal{R}_{1}:=\mathrm{M}_{n}$ and $\mathcal{R}_{2}:=\mathcal{A}$, and by taking $\pi_{\mathcal{R}}$ as the map $f:\mathcal{D}(\mathcal{R}_{2})\to T$. The assumptions of the lemma are met since $\pi_{\mathcal{A}}$ is open by Thm. 1.2.
The following two examples are concerned with openness. Continuity is a topic of Sec. 9 below.
@@
    Rem. 8.2 solves the openness problem of the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ to $\mathcal{R}$, based on the solution of the openness problem of $\pi_{\mathcal{R}}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{R})$, where $\mathcal{A}:=\mathrm{M}_{2}\oplus\mathrm{M}_{1}$. By Ex. 3.5 a), the map $\pi_{\mathcal{R}}$ is not open at any point in the half-open segment $\mathcal{G}_{0}=\big{(}\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1\big{]}\subset\mathcal{D}(\mathcal{A})$ and open at every point in the complement. By Rem. 8.2, the map $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ is not open at any point in $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{0})$ and open at every point in the complement.
@@
    $\rho(\lambda,z):=\begin{pmatrix}(1-\lambda)\ket{+}\!\!\bra{+}&z\ket{+}\\ \overline{z}\bra{+}&\lambda\end{pmatrix}\,,\quad z\in\mathbb{C}\,,\quad|z|^{2}\leq\lambda(1-\lambda)\,,$ (8.1)
@@ 9. Continuity in quantum information theory
The purpose of the maximum entropy inference method is to update a prior probability distribution if new information becomes available in the form of constraints that specify a set of possible posterior probability distributions. The preferred posterior is that which minimizes the relative entropy from the prior subject to the available constraints, see Chap. 8 in Caticha’s book [13], and the references therein. Whereas the maximum entropy method has a sound axiomatic basis, the analogous inference method for quantum states lacks an axiomatic foundation and remains a controversial topic [3].
Analytic properties of quantum mechanical inference are interesting despite unsolved foundational issues. Notably, quantum inference maps have discontinuities [55] that aroused interest in theoretical physics [14, 26, 36, 56].
@@
The relative entropy $S:\mathcal{D}(\mathrm{M}_{n})\times\mathcal{D}(\mathrm{M}_{n})\to[0,+\infty]$ is an asymmetric distance. It is defined as $S(\rho_{1},\rho_{2}):=\operatorname{tr}[\rho_{1}(\log(\rho_{1})-\log(\rho_{2}))]$ if the range of $\rho_{1}$ is included in the range of $\rho_{2}$ and as $S(\rho_{1},\rho_{2}):=+\infty$ otherwise, for all $\rho_{1},\rho_{2}\in\mathcal{D}(\mathrm{M}_{n})$. Let $\tau\in\mathcal{D}(\mathrm{M}_{n})$, the prior, be a density matrix of maximal rank $n$. Then
@@
is well defined, see [52, Def. 1.1] and the references therein.
We quote [52, Thm. 4.9] under the assumptions of Ex. 9.1.
@@
For any $\sigma\in\mathcal{D}(\mathcal{R})$, the inference map $\Psi_{\tau}:\mathcal{D}(\mathcal{R})\to\mathcal{D}(\mathrm{M}_{n})$ is continuous at $\sigma$ if and only if the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$ is open at $\Psi_{\tau}(\sigma)$.
@@
We resume Ex. 9.1. Thm. 9.2 translates the continuity problem of the inference map $\Psi_{\tau}$ into an openness problem of the orthogonal projection $\pi$. Whether Rem. 8.2 can simplify this openness problem depends on the location of the posteriors.
The range of $\Psi_{\tau}$ can be described explicitly. The manifold
$\mathcal{E}_{\tau}(\mathcal{R}):=\big{\{}\frac{e^{\log(\tau)+A}}{\operatorname{tr}(e^{\log(\tau)+A})}\mid A\in\mathcal{R}\big{\}}$ (9.2)
is known as a Gibbsian family or exponential family, see [55, 38, 52, 43] and the references therein (and recall $\mathcal{R}=\mathcal{H}(\mathcal{R})$). The manifold $\mathcal{E}_{\tau}(\mathcal{R})$ is included in $\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$. But not just that, the set $\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ of maximum entropy inference states equals
$\{\rho_{1}\in\mathcal{D}(\mathrm{M}_{n}):\inf_{\rho_{2}\in\mathcal{E}_{\tau}(\mathcal{R})}S(\rho_{1},\rho_{2})=0\}\,,$ (9.3)
which is a subset of the Euclidean closure of $\mathcal{E}_{\tau}(\mathcal{R})$, see (D5) in [52]. The set (9.3) is called reverse information closure or the rI-closure [19] of $\mathcal{E}(\mathcal{R})$.
Let $\mathcal{A}$ denote the smallest *-subalgebra $\mathcal{A}$ of $\mathrm{M}_{n}$ that contains $\mathcal{R}$.
@@
    Rem. 8.2 is of no avail if $\tau=\mathds{1}_{n}/n$ is the uniform prior. In this case,
@@
    is the von Neumann entropy $S(\rho):=-\operatorname{tr}[\rho\log(\rho)]$ up to a constant. By functional calculus, the exponential family $\mathcal{E}_{\tau}(\mathcal{R})$ is included in $\mathcal{A}$ and so is the set of inference states $\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ as per the formula (9.3), because $\mathcal{A}$ is closed. Rem. 8.2 is of no help since all posteriors lie in $\mathcal{A}$ anyways. The same conclusion is true for every prior $\tau$ in $\mathcal{A}$ replacing $\mathds{1}_{n}/n$.
@@
    If the prior $\tau$ lies outside of $\mathcal{A}$ then the exponential family $\mathcal{E}_{\tau}(\mathcal{R})$ is disjoint from $\mathcal{D}(\mathcal{A})$, again by functional calculus. Rem. 8.2 brings an advantage if the orthogonal projection $\pi_{\mathcal{R}}:\mathcal{D}(\mathcal{A})\to\mathcal{D}(\mathcal{R})$ is easier to analyze than $\pi:\mathcal{D}(\mathrm{M}_{n})\to\mathcal{D}(\mathcal{R})$. An example is given in Ex. 8.4 where $\mathcal{A}$ has a simple structure and a much smaller dimension than $\mathrm{M}_{n}$.
@@
    The real matrix system $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$ in Ex. 8.3 has the property that the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ is not open anywhere in the relative interior of the fiber over $M:=\frac{1}{2}(\ket{+}\!\!\bra{+}\oplus 1)$ and open at every point in the complement of that fiber. For every $\sigma\in\mathcal{D}(\mathcal{R})$ and prior $\tau$, the posterior $\Psi_{\tau}(\sigma)$ is contained in the relative interior of the fiber $\pi^{-1}(\sigma)$ over every $\sigma$ by Coro. 5.7 and Lemma 5.8 in [52]. Thm. 9.2 then shows that $\Psi_{\tau}$ is discontinuous at $M$ and continuous at every point different from $M$ (for all priors $\tau$).
@@
$\sum_{\nu\in\mathfrak{g}}A_{\nu}\otimes\mathds{1}_{\bar{\nu}}\,,\quad A_{\nu}\in\mathcal{H}(\mathcal{A}_{\nu})\,,\quad\nu\in\mathfrak{g}\,.$ (9.4)
@@
$\operatorname{red}_{\mathfrak{g}}:\mathcal{A}_{\Omega}\to\prod_{\nu\in\mathfrak{g}}\mathcal{A}_{\nu}\,,\quad A\mapsto[\operatorname{Tr}_{\bar{\nu}}(A)]_{\nu\in\mathfrak{g}}$ (9.5)
denote the map from $\mathcal{A}_{\Omega}$ to the cartesian product of the algebras $(\mathcal{A}_{\nu})_{\nu\in\mathfrak{g}}$ that assigns partial traces to the subsystems specified by the members of $\mathfrak{g}$.
Linear constraints on $\mathcal{D}(\mathcal{A}_{\Omega})$ have been defined in terms of reduced density matrices, see [38, 14] and [58, Sec. 1.4.2]. This is formalized in the following diagram, which commutes by formula (19) in [54]. (Obvious restrictions of $\operatorname{red}_{\mathfrak{g}}$ are omitted.)

### v2

@@ Abstract.
quantum state, matrix system, open map, numerical range
@@ 1. Introduction
So $\mathcal{O}|_{\mathcal{R}}:=\{\ell|_{\mathcal{R}}:\ell\in\mathcal{O}\}$ contains $\omega_{\lambda}|_{\mathcal{R}}$ but none of the restrictions $\ell_{\theta}|_{\mathcal{R}}$ of
@@
Corey et al. [17] and Leake et al. [29, 30] first studied a problem of numerical ranges closely related to (1.1). The problem (1.1) itself was studied by Weis [53, 54] and Rodman et al. [46] when $\mathcal{D}(\mathcal{R})$ is replaced with the affinely isomorphic joint numerical range. Numerical ranges are the topic of Sec. 4.
@@
Thm. 1.2 can simplify the problem (1.1), and related continuity problems, if $\mathcal{R}$ is included in a *-subalgebra of $\mathrm{M}_{n}$ smaller than $\mathrm{M}_{n}$ as we show in Sec. 8. Examples from quantum information theory are presented in Sec. 9.
@@ 2. States and density matrices
The aim of this section is to translate openness questions from states to density matrices. In the sequel, we refer to matrix systems and *-subalgebras synonymously as complex matrix systems and complex *-subalgebras, respectively. In addition, we introduce real *-subalgebras as they have a greater variety of state spaces than the complex ones. See, however, Rem. 2.4.
@@
Real and complex matrix systems have the same families of state spaces. If $\mathcal{R}$ is a real or complex matrix system on $\mathbb{C}^{n}$, then the real matrix system of hermitian matrices $\mathcal{H}(\mathcal{R})$ has the same set of hermitian matrices and hence the same set of density matrices as $\mathcal{R}$. Conversely, any real matrix system $\mathcal{R}$ included in $\mathcal{H}(\mathrm{M}_{n})$ has the same same set of hermitian matrices and the same set of density matrices as the complex matrix system $\mathcal{R}\oplus\mathrm{i}\mathcal{R}$. The set of density matrices of a real or complex matrix system $\mathcal{R}$ is affinely isomorphic to the state space of $\mathcal{R}$ by Lemma 2.1 or 2.2, respectively.
@@
Indeed, $\mathcal{C}_{2}^{\vee}=\mathcal{C}_{2}$ holds [28, Thm. III.2.1] as the space of hermitian matrices $\mathcal{H}_{2}$ is a Euclidean Jordan algebra with Jordan product $A\circ B=\frac{1}{2}(AB+BA)$ and inner product $(A,B)\mapsto\mathrm{Re}\langle A,B\rangle$, $A,B\in\mathcal{H}_{2}$. So, $\mathcal{C}_{2}^{\vee}=\mathcal{C}_{2}\subset\mathcal{C}_{1}\subset\mathcal{C}_{1}^{\vee}$ proves $\mathcal{D}_{2}\subset\mathcal{D}_{1}$, which implies (2.4). See also the Notes to Chapter 6 in [2].
@@ 3. Direct convex sums
As $K:=K_{1}\oplus_{\mathrm{c}}\cdots\oplus_{\mathrm{c}}K_{m}$ is a metric space, it suffices to show that $O:=O_{I}(\epsilon,(A_{i})_{i\in I})$ is open and there are arbitrary small sets of this form.
@@
    Thereby, the equivalence of states and density matrices is described in Coro. 2.7. It is important to observe that $\frac{1}{2}(\mathds{1}_{2}+\cos(\theta)X+\sin(\theta)Z)\oplus 0$ and $\cos(\theta)(X\oplus 1)+\sin(\theta)Z\oplus 0$ are matrices in $\mathrm{M}_{2}\oplus\mathrm{M}_{1}$ for all $\theta\in\mathbb{R}$. The map $\pi$ is open at every point in the complement of $\mathcal{G}_{0}$ because $\mathcal{D}(\mathrm{M}_{2}\oplus\mathrm{M}_{1})$ is a cone over a ball, see [53, Lemma 4.17] for a detailed proof.
@@
It is instructive to look at Ex. 3.5 from the perspective of a real *-subalgebra of $\mathrm{M}_{3}$, whose state space can be visualized in three-space.
@@
over $M$. The map $\pi$ is not open at any point in $\mathcal{G}_{0}=\mathcal{G}\setminus\{\ket{+}\!\!\bra{+}\oplus 0\}$ and open at every point in the complement of $\mathcal{G}_{0}$. The lack of openness can be visualized graphically in three-space by observing that $\pi$ projects the cone $\mathcal{D}(\mathrm{M}_{2}(\mathbb{R})\oplus\mathrm{M}_{1}(\mathbb{R}))$ along its generatrix $\mathcal{G}$ to the ellipse $\mathcal{D}(\mathcal{R})$.
@@ 4. Numerical ranges
$f_{A}:\mathbb{C}S^{n}\to W(A)\,,\quad\ket{\varphi}\mapsto\braket{\varphi}{A\varphi}\,.$
@@
If $z\in W(A)$ is not an extreme point of $W(A)$, then $f_{A}^{-1}$ is strongly continuous at $z$ [17, Thm. 4]. Let $z$ be an extreme point of $W(A)$. Then there are $\theta_{0}\in\mathbb{R}$ and $i_{0}\in\{1,\ldots,n\}$ such that $z=z_{i_{0}}(\theta_{0})$. Now, $f_{A}^{-1}$ is strongly continuous at $z$ if and only if $z_{i}(\theta_{0})=z$ implies $z_{i}=z_{i_{0}}$ for all $i=1,\ldots,n$, see [30, Thm. 2.1.1]. Leake at al. [30] state the latter condition by saying that the eigenfunctions corresponding to $z$ at $\theta_{0}$ do not split.
@@
Strong continuity is connected to the openness of a linear map. For all $z\in W(A)$, the multi-valued map $f_{A}^{-1}$ is strongly continuous at $z$ if and only if the restricted linear map $v:\mathcal{D}(\mathrm{M}_{n})\to V(A_{1},A_{2})$, introduced in (4.1), is open at every point in the fiber $v^{-1}[(\mathrm{Re}(z),\mathrm{Im}(z)){}^{\mathrm{T}}]$, see [54, Coro. 5.2]. Moreover, for every $x\in V(A_{1},A_{2})$, the map $v$ is open at some point in the relative interior [^2] The relative interior [45] of a convex set $K$, denoted by $\operatorname{ri}(K)$, is the interior of $K$ in the relative topology of the affine hull of $K$. of $v^{-1}(x)$ if and only if $v$ is open at every point in $v^{-1}(x)$.
@@ 5. Retractions
    The stability problem of a finite-dimensional compact convex set $K$ is completely solved [40]. The $d$-skeleton of $K$ is defined as the union of all faces [^3] A face [45] of a convex set $K$ is a convex subset $F\subset K$ such that $x,y\in F$ is implied by $(1-s)x+sy\in F$ for all $x,y\in K$ and $s\in(0,1)$. of $K$ of dimension at most $d$. The convex set $K$ is stable if and only if for every nonnegative integer $d$, the $d$-skeleton of $K$ is closed.
@@
    The set of density matrices $\mathcal{D}(\mathrm{M}_{n})$ is stable [20] because all its $d$-skeletons are closed. The closedness follows from three arguments: First, the set $\mathcal{D}(\mathrm{M}_{n})$ is a compact convex set of dimension $n^{2}-1$. Second, every nonempty face of $\mathrm{M}_{n}$ is unitarily similar to $\mathcal{D}(\mathrm{M}_{l})\oplus\{0\}$ for some positive integer $l$, and third, the unitary group $U(n)$ is compact.
@@
$K^{\times k}\to K\,,\quad(x_{1},\dots,x_{k})\mapsto\frac{1}{k}\sum_{i=1}^{k}x_{i}\,,$
defined on the $k$-fold cartesian product of $K$, is open for all positive integers $k$. More generally, for every tuple $(s_{1},\dots,s_{k})$ of nonnegative real numbers adding up to one, the map
@@ 8. Topology simplified by algebra
Indeed, the map $\pi$ factors by Lemma 2.5. The second claim follows from Lemma 8.1 a), by letting $\mathcal{R}_{1}:=\mathrm{M}_{n}$ and $\mathcal{R}_{2}:=\mathcal{A}$, and by taking $\pi_{\mathcal{R}}$ as the map $f:\mathcal{D}(\mathcal{R}_{2})\to T$. The assumptions of the lemma are met since $\pi_{\mathcal{A}}$ is open by Thm. 1.2.
The following examples demonstrate the use of Rem. 8.2. Continuity problems are a topic of Sec. 9 below.
@@
Keeping the notation of $\pi=\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$ from Rem. 8.2, we take $\mathcal{R}:=\operatorname{span}_{\mathbb{R}}(\mathds{1}_{3},X\oplus 1,Z\oplus 0)$ and $\mathcal{A}:=\mathrm{M}_{2}\oplus\mathrm{M}_{1}$.
@@
    By Ex. 3.5 a), the map $\pi_{\mathcal{R}}$ is not open at any point in the half-open segment $\mathcal{G}_{0}=\big{(}\ket{+}\!\!\bra{+}\oplus 0,0\oplus 1\big{]}\subset\mathcal{D}(\mathcal{A})$ and open at every point in the complement. By Rem. 8.2, the map $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ is not open at any point in $\pi_{\mathcal{A}}^{-1}(\mathcal{G}_{0})$ and open at every point in the complement.
@@
    $\rho(\lambda,z):=\begin{pmatrix}(1-\lambda)\ket{+}\!\!\bra{+}&z\ket{+}\\ \overline{z}\bra{+}&\lambda\end{pmatrix}\,,\quad z\in\mathbb{C}\,,\quad|z|^{2}\leq\lambda(1-\lambda)\,,$
@@ 9. Continuity in quantum information theory
The purpose of the maximum entropy inference method is to update a prior probability distribution if new information becomes available in the form of constraints that specify a set of possible posterior probability distributions. The preferred posterior is that which minimizes the relative entropy from the prior subject to the available constraints, see Chap. 8 in Caticha’s book [12] and the references therein. An analogous quantum mechanical inference method can be defined by replacing probability distributions with density matrices and the standard relative entropy with the Umegaki relative entropy. The axiomatic foundations of the maximum entropy inference method were settled for probability distributions in the 1980’s, see Chap. 6 in [12, pp. 157–160]. More than 30 years later, the search for axioms of the quantum inference [3] found a convincing answer in the work of Vanslette [51].
@@
The relative entropy $S:\mathcal{D}(\mathrm{M}_{n})\times\mathcal{D}(\mathrm{M}_{n})\to[0,+\infty]$ is an asymmetric distance. It is defined by $S(\rho_{1},\rho_{2}):=\operatorname{tr}[\rho_{1}(\log(\rho_{1})-\log(\rho_{2}))]$ if the range of $\rho_{1}$ is included in the range of $\rho_{2}$ and by $S(\rho_{1},\rho_{2}):=+\infty$ otherwise, for all $\rho_{1},\rho_{2}\in\mathcal{D}(\mathrm{M}_{n})$. Let $\tau\in\mathcal{D}(\mathrm{M}_{n})$, the prior, be a density matrix of maximal rank $n$. Then
@@
is well defined, see [53, Def. 1.1] and the references therein. Discontinuities of this inference map [56] aroused interest in theoretical physics [13, 26, 36, 57]. The map (9.1) is continuous for commutative real matrix systems $\mathcal{R}$, for example for the inference of probability distributions mentioned above.
We quote [53, Thm. 4.9].
@@
Let $\sigma\in\mathcal{D}(\mathcal{R})$. Then $\Psi_{\tau}$ is continuous at $\sigma$ if and only if $\pi$ is open at $\Psi_{\tau}(\sigma)$.
In what follows, continuity and openness problems will be simplified by factoring $\pi=\pi_{\mathcal{R}}\circ\pi_{\mathcal{A}}$ through a *-subalgebra $\mathcal{A}$ of $\mathrm{M}_{n}$ that includes $\mathcal{R}$, using the notation of Rem. 8.2.
@@
Before simplifying the openness problem of $\pi$, we describe the set $\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ of posteriors, as the openness only matters for points in this set by Thm. 9.2. Recalling $\mathcal{R}=\mathcal{H}(\mathcal{R})$, we define
$\mathcal{E}_{\tau}(\mathcal{R}):=\big{\{}\frac{e^{\log(\tau)+A}}{\operatorname{tr}(e^{\log(\tau)+A})}\mid A\in\mathcal{R}\big{\}}\,.$
The manifold $\mathcal{E}_{\tau}(\mathcal{R})$ is known as a Gibbsian family or exponential family, see [56, 38, 53, 43] and the references therein. By (D5) in [53] we have
$\Psi_{\tau}(\mathcal{D}(\mathcal{R}))=\{\rho_{1}\in\mathcal{D}(\mathrm{M}_{n}):\inf_{\rho_{2}\in\mathcal{E}_{\tau}(\mathcal{R})}S(\rho_{1},\rho_{2})=0\}\,.$ (9.2)
The right-hand side of (9.2) is the reverse information closure or rI-closure [18] of $\mathcal{E}_{\tau}(\mathcal{R})$, which is a subset of the Euclidean closure of $\mathcal{E}_{\tau}(\mathcal{R})$.
@@
    If $\tau=\mathds{1}_{n}/n$ is the uniform prior, then
@@
    is the von Neumann entropy $S(\rho):=-\operatorname{tr}[\rho\log(\rho)]$ up to a constant. By functional calculus, $\mathcal{E}_{\tau}(\mathcal{R})$ is included in $\mathcal{A}$ and so is the set of posteriors $\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ as per (9.2), because $\mathcal{A}$ is closed. Rem. 8.2 then shows that for every $\rho\in\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ the openness of $\pi$ at $\rho$ is equivalent to the openness of $\pi_{\mathcal{R}}=\pi|_{\mathcal{D}(\mathcal{A})}$ at $\rho$ (the same conclusion is true for every prior $\tau$ in $\mathcal{A}$). This can simplify the problem if $\mathcal{A}$ has a simpler structure or a smaller dimension than $\mathrm{M}_{n}$. An example is given in Ex. 8.4 above.
@@
    If the prior $\tau$ lies outside of $\mathcal{A}$ then $\mathcal{E}_{\tau}(\mathcal{R})$ is disjoint from $\mathcal{D}(\mathcal{A})$, again by functional calculus. Rem. 8.2 then shows that for every $\rho\in\Psi_{\tau}(\mathcal{D}(\mathcal{R}))$ the openness of $\pi$ at $\rho$ is equivalent to the openness of $\pi_{\mathcal{R}}$ at $\pi_{\mathcal{A}}(\rho)$. This is an even greater simplification than in part a) above, because the analysis of $\pi$ on $\mathcal{D}(\mathrm{M}_{n})$ is reduced to that of $\pi_{\mathcal{R}}$ on $\mathcal{D}(\mathcal{A})$.
@@
    Independently of $\mathcal{A}$, it sometimes helps that the posterior $\Psi_{\tau}(\sigma)$ is contained in the relative interior of the fiber $\pi^{-1}(\sigma)$ over $\sigma$ for every $\sigma\in\mathcal{D}(\mathcal{R})$ and prior $\tau$ by Coro. 5.7 and Lemma 5.8 in [53]. As an example, the orthogonal projection $\pi:\mathcal{D}(\mathrm{M}_{3})\to\mathcal{D}(\mathcal{R})$ to the real matrix system $\mathcal{R}$ in Ex. 8.3 is not open anywhere in the relative interior of the fiber over a certain point $M$ and open at every point in the complement of that fiber. Thm. 9.2 then shows that $\Psi_{\tau}$ is discontinuous at $M$ and continuous everywhere else in the ellipse $\mathcal{D}(\mathcal{R})$ for every prior $\tau$.
From here on, we assume the prior $\tau:=\mathds{1}_{n}/n$ be uniform. Using the von Neumann entropy $S=\phi_{\tau}+\log(n)$, we write the inference map (9.1) from Ex. 9.1 as
We also write $\mathcal{E}(\mathcal{R}):=\mathcal{E}_{\tau}(\mathcal{R})$ for the exponential family of Ex. 9.3.
@@
$\sum_{\nu\in\mathfrak{g}}A_{\nu}\otimes\mathds{1}_{\bar{\nu}}\,,\quad A_{\nu}\in\mathcal{H}(\mathcal{A}_{\nu})\,,\quad\nu\in\mathfrak{g}\,.$
@@
$\operatorname{red}_{\mathfrak{g}}:\mathcal{A}_{\Omega}\to\prod_{\nu\in\mathfrak{g}}\mathcal{A}_{\nu}\,,\quad A\mapsto[\operatorname{tr}_{\bar{\nu}}(A)]_{\nu\in\mathfrak{g}}$
denote the map from $\mathcal{A}_{\Omega}$ to the cartesian product of the algebras $(\mathcal{A}_{\nu})_{\nu\in\mathfrak{g}}$ that assigns reduced density matrices.
Linear constraints on $\mathcal{D}(\mathcal{A}_{\Omega})$ have been defined in terms of reduced density matrices, see [38, 13] and [59, Sec. 1.4.2]. This is formalized in the following diagram, which commutes by formula (19) in [55]. (Obvious restrictions of the domain and codomain of $\operatorname{red}_{\mathfrak{g}}$ are omitted in the sequel.)